OpenStudy (anonymous):

Differentiate:\[-\frac{d}{dx}f(x) = \frac{d}{dx}f(-x)\]I am trying to prove that the derivative of an odd function is even, but when I try to use the Chain Rule, it comes out to the derivative = 1

OpenStudy (anonymous):

Did I miss a negative symbol somewhere?

OpenStudy (anonymous):

maybe I misunderstand the problem, but if you derive the lefthand side you obtain the negative derivative of the general function right? \[ -f'(x)\] ?

OpenStudy (anonymous):

The question is just: Given that f(x) is odd, prove that f'(x) = f'(-x)

OpenStudy (anonymous):

well for the right hand side I would use the chain rule, then you will get \[ -f'(x)=-f'(-x) \] the negative cancels out.

OpenStudy (anonymous):

(in fact I used the chain rule for both of them, but on the LHS it's just multiplying by 1) on the right hand side it's multiplying by -1

OpenStudy (anonymous):

Are you sure it's not f(x)=-f(-x) to differentiate?

OpenStudy (anonymous):

wouldn't make a difference, if you differentiate what you just have written you get the same

OpenStudy (anonymous):

\[\frac{d}{dx}f(x) = -\frac{d}{dx}f(-x)\]right?

OpenStudy (anonymous):

I know the definition of an odd function as the following: \[ -f(x)=f(-x)\]

OpenStudy (anonymous):

Yes, I just moved the negative sign over

OpenStudy (anonymous):

good, so we're doing the same, yes, now differentiate, using the chain rule.

OpenStudy (anonymous):

\[\frac{d}{dx}f(x) = -\frac{d}{dx}f(-x) \cdot \frac{d}{dx}f(x)\]am I applying the Chain Rule correctly?

OpenStudy (anonymous):

The chain rule says, differentiate the outside with respect to the inside and then differentiate the inside and multiply, so what you write above doesn't look bad, but I believe this is more what happens: \[ \Large \frac{d}{dx} f(x)=-\frac{d}{dx}f(-x)\cdot \frac{d}{dx}(-x)\]

OpenStudy (anonymous):

Whoa, okay, I guess I don't fully understand the Chain Rule yet. I have no idea why you applied\[\frac{d}{dx}(-x)\]for the Chain Rule

OpenStudy (anonymous):

\[ \frac{d}{dx}(5x+7)^8 =8(5x+8)^7 \cdot 5 \] First I derived the outside of the function with respect to the inside, hence the inside stays the same, notice on both LHS and RHS we have 5x+7, but the function (the exponential) got derived. 8f(x)^7 and then I multiplied times the derivative of the inside which is 5.

OpenStudy (anonymous):

oh sorry, I misspelled, it should be 5x+7 on both of course, my typing is bad today.

OpenStudy (anonymous):

Ohh, I see now. I was using the Chain Rule similarly as I would with implicit differentiation

OpenStudy (anonymous):

CORRECTION ! : \[ \large \frac{d}{dx}(5x+7)^8 =8(5x+7)^7 \cdot 5\]

OpenStudy (anonymous):

So we have\[\frac{d}{dx}f(x) = -\frac{d}{dx}f(-x) \cdot \frac{d}{dx}(-x)\]Not sure how to combine these terms or what I'm allowed to do and not do

OpenStudy (anonymous):

The derivative of f(x) is just f'(x), similarly for f(-x) the derivative is f'(-x) and the derivative of x would be one as you know, but you differentiate -x here, so the derivative becomes minus one.

OpenStudy (anonymous):

therefore the minus cancels out, as predicted by the comment above we had to prove.

OpenStudy (anonymous):

Usually I see derivatives in the form:\[\frac{dy}{dx}\]so this kind of form really threw me off...is this the same as\[\frac{d}{dx}y\]?

OpenStudy (anonymous):

yes this is Leibniz's Notation for derivatives, it's a bit confusing at first, the term \[ \large \frac{d}{dx} \] is Leibniz differential Operator, there are many equal ways of writing something here now: \[ \large f'(x)= \frac{dy}{dx}=\frac{d}{dx}y = \frac{d(f(x))}{dx} \]

OpenStudy (anonymous):

you will see plenty of those.

OpenStudy (anonymous):

So isn't\[\frac{dy}{dx} = 1\]?

OpenStudy (anonymous):

No why should it be? The derivative of a single variable is one, if you differentiate it to respect of that variable \[ y= x \] Apply Leibniz Operator \[ \frac{dy}{dx}=\frac{dx}{dx}=1 \]

OpenStudy (anonymous):

y is a function, it can be anything. From linear to complex.

OpenStudy (anonymous):

If f(x) = y How would I write f(-x) in terms of y?

OpenStudy (anonymous):

Same, you don't have a function specified there, remember that whatever you denote into the brackets you will write in the actual function \[ f(-x)=y\] this is an even function now

OpenStudy (anonymous):

But you can write -f(x) as -y, right?

OpenStudy (anonymous):

yes you can do that \[ -y = f(-x) \\ y=-f(-x)\] Is this what you are trying to do? Differentiate \[ \frac{dy}{dx}=f'(x)=f'(-x)\]

OpenStudy (anonymous):

Something like that, I'm just trying to understand the general concepts

OpenStudy (anonymous):

same result as above, so it works